OpenStudy (anonymous):

What is the slope of the line passing through the point (4, -2) and perpendicular to the line y=4x+2? Could possibly be written as a fraction... I think.

ganeshie8 (ganeshie8):

start by finding the slope of "given" line

ganeshie8 (ganeshie8):

whats the slope of line "y = 4x+2" ?

OpenStudy (anonymous):

Would the slope of the line be 4?

ganeshie8 (ganeshie8):

yes! next use below to calculate the slope of perpendicular line :-

ganeshie8 (ganeshie8):

if two lines are perpendicular, then product of slopes = -1

ganeshie8 (ganeshie8):

say, the requires slope is \(m\) :- \(m \times 4 = -1\)

ganeshie8 (ganeshie8):

that gives \(\large m = \frac{-1}{4}\)

ganeshie8 (ganeshie8):

so the perpendicular line will have a slope of \(\large \frac{-1}{4}\)

ganeshie8 (ganeshie8):

if that makes some sense.. .

OpenStudy (anonymous):

Oh wow, much easier to understand than I thought it would be.

OpenStudy (anonymous):

So its -1/4? I wasn't sure if it would be a fraction.

ganeshie8 (ganeshie8):

yes, -1/4 is the required slope :)

OpenStudy (anonymous):

Thank you very much.

ganeshie8 (ganeshie8):

np.. u wlc :)